OpenStudy (anonymous):

What is the sum of the geometric sequence –1, 6, –36, … if there are 6 terms?

OpenStudy (anonymous):

S6 = a * (r^6 - 1) / ( r - 1) a = -1 and r = -36/6 = -6 so Sum 6 terma = -1* (-6 ^6 -1 ) / ( -6-1) = 6665